# Hierarchical Mixture of Experts: Generalizable Learning for High-Level Synthesis

### v1

@@ Appendix D Supplementary Experiment Results
As analyzed in the Experiments section in the main paper, one concern is whether the hierarchical MoE model simply benefits from having more parameters. To address this concern, we enlarge the hidden size of HARP from 64 to 128 and 256, and we evaluate their performance. The results are listed in Table 10. As we enlarge the hidden size, the MSE also increases. Even if the parameter size is larger than hierarchical MoE, its performance is still worse than hierarchical MoE. In the original best hyper-parameter setting of HARP, 64 is already the best-performing hidden size. We also list the parameter size of single-level MoE models in Table 11. If we do not utilize the hierarchical structure, only using MoE on one granularity does not significantly increase the parameter size. It verifies that parameter size is not the key reason for MoE’s success.
@@
Another concern is about the number of experts. In the hierarchical MoE, if we only utilize MoE on the node granularity or graph granularity, its performance is slightly worse than utilizing MoE on three granularities, but it could still perform well (Table 4 in the main paper). In this case, the three expert models for the high-level MoE have the same structure. Each of the low-level MoE models has 4 experts, so in total we have 12 experts for that granularity and 3 experts for other layers that are not equipped with low-level MoE. Does its success come from a large number of expert models (12 experts for that granularity), or the hierarchical structure (12 experts for that granularity + 3 experts for other layers)? To answer this question, we run experiments on the single-level MoE structure with different numbers of experts. In addition to our original setting using 4 experts, we experiment with 2, 8, and 12 experts. The results are listed in Table 12. Increasing the number of experts does not necessarily improve the performance. For all three MoEs, the lowest MSE is achieved when we use 4 experts. Therefore, the performance gain is probably because of the hierarchical structure (12 experts for that granularity and 3 experts for other layers).
@@
Due to the limited space, we only analyze the gating network of the high-level MoE and the block MoE in the main paper. Here we analyze the gating network of the node MoE and the graph MoE. We still pick the best-performing hierarchical MoE model from the five repeated experiments to analyze. For the node MoE, we show the averaged assigned weights for each type of node in Table 14. There are three non-overlapping types of nodes: normal nodes that represent the data and instructions, pseudo nodes that represent basic blocks, and pragma nodes that represent pragmas. A special type of normal node is icmp node, which represents the “icmp” instruction of a loop. If a pragma modifies a loop, that pragma node will be connected to both the icmp node and the pseudo node corresponding to the loop, so the icmp nodes are very meaningful. The normal nodes and pseudo nodes are relatively equally distributed to the experts. The icmp nodes are mainly handled by the third expert, and the pragma nodes are mainly handled by the fourth expert. Similar to our analysis of the block MoE in the main paper, the node MoE’s experts also diversify their roles.
@@
Table 16: The scope of MoE-related hyper-parameters that we have tried. “Choice” is our final choice.
@@
Node MoE’s place every layer in the GNN encoder, the GNN layer after pragma MLP*, additional linear layer marked
Block MoE’s place pragma MLP, pooling, additional linear layer, additional linear layer and remove some layers* marked
For graph MoE, based on the expert assignment weights of each data point, we calculate the average expert assignment weights of each kernel. Then, we can calculate the cosine similarity between two kernels using the weights. Table 15 lists the most similar source kernel of each target kernel. “Gemver-medium” and its most similar kernel, “gesummv-medium”, are both matrix-vector multiplications; “syr2k” and its most similar kernel, “syrk’, are both symmetric rank operations; “trmm-opt” and its most similar kernel, “symm”, are both matrix multiplications. This indicates that the expert weights assigned by the graph MoE are partially explainable.
@@ Appendix E Other Detailed Information
Hyper-parameters. We basically follow the hyper-parameter settings of HARP. We use the Adam optimizer to train the model for 1000 epochs and fine-tune it for 500 epochs. The learning rate is 1e-3 and the weight decay is 1e-4. We use the cosine annealing learning rate schedule (Loshchilov and Hutter 2017) with a minimum learning rate of 1e-5, and the unturned linear warmup (Ma and Yarats 2021). We set the hidden size as 64 and the dropout as 0.1. The regression models use the MSE loss to train, while the classification models use the cross entropy loss. In the first design of the high-level gating network, we use the graph pooling of input features (Equation 5 in the main paper). Its MLP has two layers and one ReLU function between them, where the first layer has the same hidden size as the input dimension and the second one shrinks the output dimension to 1.
For hyper-parameters related to MoE, we try the hyper-parameters in a reasonable scope and use the best one with the lowest test MSE on the target kernels. The low-level MoE uses the weighted sum of 4 experts, and the high-level MoE calculates the weighted sum of three models (the node MoE model, the block MoE model, and the graph MoE model). The gating network is a linear layer and a softmax function. We set the weight of the regularization term of MoE, both $\alpha$ and $\beta$, as 5e-3. In the two-stage training of the high-level MoE, the first stage (warmup) takes the first 500 epochs, and the second stage takes the later 500 epochs. In the model variants experiments, we only modify the different parts as introduced in the experiments. Table 16 lists the scope of hyper-parameters we have tried.
Reproducibility. We have submitted our data and codes in the supplementary material. The readme file contains instructions on how to run the codes. We have introduced the important hyper-parameters in this appendix, while other hyper-parameters can be found in the “src/config.py”, We use NVIDIA L40S (40GB) to train our model. It takes about 5 hours to train the model on the source kernels and about 1 hour to fine-tune on the target kernels. Training costs about 30 GB of memory and fine-tuning costs about 12 GB of memory. Now we load the whole dataset into the memory before training. If the memory is not sufficient, we have an argument to disable this feature conveniently, and the memory cost for training can be reduced to about 15 GB. We run HLS on an AMD EPYC 7V13 64-core processor. We run HLS for the selected top-10 design points in parallel. It takes about 10 CPUs, 50 GB of memory, and 3 hours to run 10 design points for one kernel.
@@ Appendix F Limitations and Future Work
Though we have made much progress on domain generalization for HLS prediction, we also face many limitations. First, the hierarchical MoE model does not have a huge advantage over HARP on every kernel. As shown in Table 1 in the main paper, it is significantly better than HARP on “Fd” and “Gemv”, slightly better on “Gemm” and “Tr”, but it is the same as HARP or worse than HARP on “Sy” and “Ja”. This might be due to the inherent difficulty of this problem. Future research is needed to develop even stronger models.
Besides, the hierarchical MoE structure is mainly useful in quick domain adaption using a small amount of data to fine-tune. However, when the computing resource is abundant enough to run a lot of HLS on the target kernels, we will have more data points for tine-tuning, and in this case, MoE would perform similarly to HARP. The proposed model should be used in suitable applications. It also remains an open question of how to enhance the model’s general ability when the data is abundant.

### v2

@@ Appendix D Supplementary Experiment Results
One concern is whether the hierarchical MoE model simply benefits from having more parameters. To address this concern, we enlarge the hidden size of HARP from 64 to 128 and 256, and we evaluate their performance. The results are listed in Table 11. As we enlarge the hidden size, the MSE increases, so its performance becomes worse. In the original best hyper-parameter setting of HARP, 64 is already the best-performing hidden size. We also list the parameter size of single-level MoE models in Table 12. If we do not utilize the hierarchical structure, only using MoE on one granularity does not significantly increase the parameter size, but it decreases the MSE. It verifies that parameter size is not the key reason for MoE’s success.
@@
Another concern is about the number of experts. In the hierarchical MoE, if we only utilize MoE on the node granularity or graph granularity, its performance is slightly worse than utilizing MoE on three granularities, but it could still perform well (Table 5 in the main paper). In this case, the three expert models of the high-level MoE have the same structure. Each of the low-level MoE models has 4 experts, so in total we have 12 experts for that granularity and 3 experts for other layers that are not equipped with low-level MoE. Does its success come from a large number of expert models (12 experts for that granularity), or the hierarchical structure (12 experts for that granularity + 3 experts for other layers)? To answer this question, we run experiments on the single-level MoE structure with different numbers of experts. In addition to our original setting using 4 experts, we experiment with 2, 8, and 12 experts. The results are listed in Table 13. The results demonstrate that increasing the number of experts does not necessarily improve the performance. For all three MoEs, the lowest MSE is achieved when we use 4 experts. Therefore, the performance gain is probably because of the hierarchical structure (12 experts for that granularity and 3 experts for other layers).
@@
It is interesting to see whether the expert assignment is meaningful. Due to the limited space, we only analyze the gating network of the high-level MoE and the block MoE in the main paper. Here we analyze the gating network of the node MoE and the graph MoE. We still pick the best-performing hierarchical MoE model from the five repeated experiments to analyze. For the node MoE, we show the averaged assigned weights for each type of node in Table 15. There are three non-overlapping types of nodes: normal nodes that represent the data and instructions, pseudo nodes that represent basic blocks, and pragma nodes that represent pragmas. A special type of normal node is icmp node, which represents the “icmp” instruction of a loop. If a pragma modifies a loop, that pragma node will be connected to both the icmp node and the pseudo node corresponding to the loop, so the icmp nodes are very meaningful. The normal nodes and pseudo nodes are relatively equally distributed to the experts. The icmp nodes are mainly handled by the third expert, and the pragma nodes are mainly handled by the fourth expert. Similar to our analysis of the block MoE in the main paper, the node MoE’s experts also diversify their roles.
@@
Table 17: The scope of MoE-related hyper-parameters that we have tried. “Choice” is the best choice (our final choice).
@@
Node MoE’s place every GNN layer in the GNN encoder, the GNN layer after pragma MLP*, additional linear layer *
Block MoE’s place pragma MLP, pooling, additional linear layer, additional linear layer and remove some layer* *
For graph MoE, based on the expert assignment weights of each data point, we calculate the average expert assignment weights of each kernel. Then, we can calculate the cosine similarity between two kernels using the weights. Table 16 lists the most similar source kernel for each target kernel. “Gemver-medium” and its most similar kernel, “gesummv-medium”, are both matrix-vector multiplications; “syr2k” and its most similar kernel, “syrk’, are both symmetric rank operations; “trmm-opt” and its most similar kernel, “symm”, are both matrix multiplications. This indicates that the expert weights assigned by the graph MoE are partially explainable.
@@ Appendix E Other Detailed Information
Hyper-parameters. We basically follow the hyper-parameter settings of HARP. We use the Adam optimizer to train the model for 1000 epochs and fine-tune it for 500 epochs. The learning rate is 5e-4 and the weight decay is 1e-4. We use the cosine annealing learning rate schedule (Loshchilov and Hutter 2017) with a minimum learning rate of 1e-5, and the unturned linear warmup (Ma and Yarats 2021). We set the hidden size as 64 and the dropout as 0.1. The regression models use the MSE loss to train, while the classification models use the cross entropy loss. In the first design of the high-level gating network, we use the graph pooling of input features (Equation 5 in the main paper). Its MLP has two layers and one ReLU function between them, where the first layer has the same hidden size as the input dimension and the second one shrinks the output dimension to 1.
For hyper-parameters related to MoE, we try the hyper-parameters in a reasonable scope and use the best one with the lowest training-set MSE on the target kernels. The low-level MoE uses the weighted sum of 4 experts, and the high-level MoE calculates the weighted sum of three models (the node MoE model, the block MoE model, and the graph MoE model). The gating network is a linear layer and a softmax function. We set the weight of the regularization term of MoE, both $\alpha$ and $\beta$, as 5e-3. In the two-stage training of the high-level MoE, the first stage (warmup) takes the first 500 epochs, and the second stage takes the later 500 epochs. In the model variants’ experiments, we only modify the changing hyper-parameters as introduced in the experiments. Table 17 lists the scope of hyper-parameters we have tried.
Reproducibility. We have publicized our codes at https://github.com/weikai-li/HierarchicalMoE. The readme file contains instructions on how to run the codes. We have introduced the important hyper-parameters in this appendix, while other hyper-parameters can be found in the “src/config.py”, We use NVIDIA L40S (40GB) to train our model. It takes about 5 hours to train the model on the source kernels and about 1 hour to fine-tune it on the target kernels. Training costs about 30 GB of memory and fine-tuning costs about 12 GB of memory. We load the whole dataset into the memory before training. If the memory is not sufficient, we have an argument to disable this feature conveniently, and the memory cost for training can be reduced to about 15 GB. We run HLS on an AMD EPYC 7V13 64-core processor. We run HLS for the selected top-10 design points in parallel. It takes about 10 CPUs, 50 GB of memory, and 3 hours to run 10 design points for one kernel.
@@ Appendix F Limitations and Future Work
Though we have made much progress on domain generalization for HLS prediction, domain generalization is still a big challenge for HLS prediction models. First, the hierarchical MoE model does not have a huge advantage over HARP on every kernel. As shown in Table 1 in the main paper, it is significantly better than HARP on “Fd” and “Gemv”, slightly better on “Gemm” and “Tr”, but it is the same as HARP or worse than HARP on “Sy” and “Ja”. This might be due to the inherent difficulty of this problem. Future research is needed to develop even stronger models.
Second, HLS prediction is very difficult on complex kernels. In our experiments of domain transfer on complex kernels, the MSE is much larger than on the HLSyn kernels, as shown in Table 3. Although hierarchical MoE performs better than the baseline models, its MSE is still quite large. It demonstrates that hierarchical MoE is useful in improving the domain generalization ability, but it remains a future work to generally improve the model’s ability of HLS prediction on complex kernels.
Besides, the hierarchical MoE structure is mainly useful in quick domain adaption using a small amount of data to fine-tune. When the computing resource is abundant enough to run a lot of HLS on the target kernels, we will have more data points for tine-tuning, and in this case, the hierarchical MoE model would perform similarly to HARP. The proposed model should be used in suitable applications.